determine whether (-6,-8) is a solution of 5x-5y=6
Answer by AAfter.Com(18)

For (-6,-8), L.H.S. of the equality sign of the equation 5x-5y=6 becomes 5x(-6) - 5x(-8) = -30 + 40 = 10
R.H.S. of the equality sign is 6
Hence, L.H.S. is not equal to R.H.S.
Therefore, (-6,-8) is not a solution of the given equation.